OpenStudy (anonymous):

The points (3, -10) and (6, r) lie on a line with the slope 2. Find the missing coordinate r.

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{r--10=?}{6-3=?}=slope\] \[\LARGE \frac{r+10=?}{6-3=?}=slope\]

pooja195 (pooja195):

What is 6-3=?

pooja195 (pooja195):

@answersforquestions are you there?? We can work on this and I know we can find the answer! :)